MINUTES FROM THE BOOSTER SUPERVISOR MEETING
Links: Minutes, PSB home page, PSB
PRESENT: M. Benedikt, A. Blas, C. Carli, G. Cyvoct, G. Graewer, C.
Jensen, K. Hanke, M. Lindroos, J. Sanchez, R. Scrivens.
Supervisor Linac: G. Graewer à K. Hanke
Supervisor PSB: A. Blas à M. Benedikt
- Linac 2: Tuesday: ignitron problem caused sparking on the
delay line for the RF of tank1 and RFQ.
- Tail clipper timing problems on Wednesday: A missing 50 Ohm termination
on pulse repeaters stopped the chain working faster than 600 ms. When
the tail clipper timing LX.TCL-CPS was added with its release trigger, it
blocked all tail clippers. (The pulse repeater was cabled 3 years ago!!!).
- On Monday, the change of a fuse in the Linac water station accidentally
switched off the station. It took some time to restart the RF tank1.
- Linac 3: lead ion operation.
- Tuesday, a regulation card on the ejection septum BE.SMH15L1 had to be
changed, hopefully this ends the series of problems that we had with the
- Saturday, the ejection kicker ring 2 BE2.KFA14L1 went down, a
microprocessor card was taken out and put back in which solved the problem.
- ISOLDE radiation alarm: during the weekend there was a radiation alarm
detected in the ISOLDE Zone HRS (Monitor PAXY05). A warning is given for
radiation levels above 270, alarm at levels above 300 mikroSv. As a
consequence the Booster operator reduced the beam intensity. However, in
such a case the ISOLDE engineer in charge should be called as this type of
problem is not directly related to the Booster beam.
- Booster - PS synchronization: The synchronization references for
both machines for all the combinations of different beam control systems
have been established and are available on the PS/PSB web pages and as paper
- We had problems with the BTU pick-ups for the SFTPRO beam caused by a
false information on the harmonic number of the beam (was set to h=1 instead
of h=2). The "knob" containing this info is in the working set
"RF global control" and serves only for this information on the
telegram. The rf-team will try to find a more significant indication for the
- Status of the different beams: No specific problems
- MDs: Studies on LHC pilot beam have started (User MDSPS).
- Next BS meeting Tuesday 4-June-02 at 10.00 in the MCR Glassbox.
Minutes by M. Benedikt.